Lesson 14 7.Aluminum oxide is decomposed using electricity to produce aluminum metal. How many grams of aluminum metal can be produced from 100.0 g of Al2O3?

2Al2O3 = 4Al + 3O2

8.Predict the mass of hydrogen gas produced when 2.73 g of aluminum reacts in a single replacement reaction with excess sulphuric acid.

2Al + 3H2SO4 = 3H2 + Al2(SO4)3


1
Expert's answer
2021-03-23T04:03:38-0400

1. 2Al2O3 = 4Al + 3O2

m(Al) = [m(Al2O3) × 4 × Mr(Al)] / [2 × Mr(Al2O3)] = [100.0 g × 4 × 26.98 g/mol] / [2 × 101.96] = 52.92 g


2. 2Al + 3H2SO4 = 3H2 + Al2(SO4)3

m(H2) = [m(Al) × 3 × Mr(H2)] / [2 × Mr(Al)] = [2.73 g × 3 × 2] / [2 × 26.98 g/mol] = 0.304 g


Answer: 52.92 g; 0.304 g